40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri May 2 03:51:29 2014, rana, Update, LSC, farther into CM MCkicked.png
    Reply  Sun May 4 22:28:54 2014, ericq, Update, LSC, farther into CM OLTF.png
       Reply  Tue May 6 03:17:15 2014, rana, Update, LSC, farther into CM 
          Reply  Tue May 6 17:58:44 2014, ericq, Update, LSC, farther into CM AOTFs.pdf
Message ID: 9899     Entry time: Fri May 2 03:51:29 2014     Reply to this: 9908
Author: rana 
Type: Update 
Category: LSC 
Subject: farther into CM 

Rana, Q

After some more matlab loopology (see Qlog), we turned on the AO path successfully. The key was to turn on the 300:80 filter in the MCL path so that it could cross stably with the AO. Then we ramp up the AO gain via the newly AC coupled AO path into the MC servo board.

The POY11 signal looks nice and smooth. For the final smoothness after the overall common gain is ramped up, I turned on a FM7 pole at 300 Hz so that the MC path would keep falling like 1/f^2 and not interfere with the AO path around 1 kHz.

There's not enough gain yet to be able to turn on the Boost. PCDRIVE is ~3 V. Earlier tonight we were seeing the EOM saturation effect maybe, but we re-allocated the gain more to the front end and its all fine now. I think we can get another ~10-15 dB of gain by using the POY whitening gain slider + the CM AO slider. Then we can get the Boost on and take some TFs with the SR785 (as long GPIB allows).

Good Settings:

CM REFL1 = +31 dB, AO = +16 dB, MC IN2 = +16 dB. SUS-MC2_LSC = FM6, FM& ON

 

** Everything has been pretty stable tonight except some occasional MC/EOM locking oscillations. This means that its been easy to keep trying some different CM steps since the Y-Arm relocks using MCL within seconds.

Attachment 1: MCkicked.png  608 kB  Uploaded Fri May 2 04:54:38 2014  | Hide | Hide all
MCkicked.png
ELOG V3.1.3-